Suppose a LMC manufactures Wild Hyacinth, a perfume, which sells them for $50. The annual costs incurred in the production and sale of Wild Hyacinth are $200,000. Each bottle of Wild Hyacinth requires $10 in materials, labor, and energy.

Consider the previous problem. Another company, TTS, has offered to produce Wild Hyacinth for LMC for $22 a bottle.

  

  1. At which quantity is it more economical for LMC to make versus buy?

  1. Given the LMC’s desire to make $100,000 in profit annually, should they make or buy the product?
Answer #1

1)

Quantity at which it is more economical for LMC to make than buy = Annual cost incurred in production and sale / (Variable cost of buy - Variable production cost)

= 200000/(22 - 10)

= 16,667 bottles

-----------------

2)

Volume required to manufacture to make $100,000 in profit annually

= (200000+100000)/(50-10)

= 7500 bottles

Volume required to Buy to make $100,000 in profit annually

= 100000/(50-22)

= 3572 bottles

Less volume is required for Buy option .

Therefore they should buy